40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Dec 14 19:29:42 2021, Paco, Update, BHD, 1Y0 rack work for LO1 
    Reply  Mon Jan 3 12:05:08 2022, Paco, Update, BHD, 1Y0 rack work for LO2 AS1 AS4 
       Reply  Mon Jan 3 16:46:41 2022, Paco, Update, BHD, 1Y1 rack work for SR2, PR2, PR3 
          Reply  Fri Jan 7 17:54:13 2022, Anchal, Update, BHD, PR2 Sat Amp has a bad channel Screen_Shot_2022-01-07_at_6.05.39_PM.png
             Reply  Fri Jan 7 18:28:13 2022, Koji, Update, BHD, PR2 Sat Amp has a bad channel 
                Reply  Mon Jan 10 13:35:52 2022, Anchal, Update, BHD, PR2 Sat Amp has a bad channel 
                   Reply  Mon Jan 10 15:59:46 2022, Koji, Update, BHD, PR2 Sat Amp has a bad channel 6x
                      Reply  Mon Jan 10 17:04:47 2022, Anchal, Update, BHD, AS1 Sat Amp CH2 had offset 
Message ID: 16540     Entry time: Mon Jan 3 16:46:41 2022     In reply to: 16539     Reply to this: 16555
Author: Paco 
Type: Update 
Category: BHD 
Subject: 1Y1 rack work for SR2, PR2, PR3 

[Paco, Anchal]

Continued working on 1Y1 rack. Populated the 6 coil drivers, made all connections between sat amp, AA chassis, DAC, and ADC adapters for SR2, PR2, and PR3 suspensions. Powered all boxes and labeled them and cables where needed. Near the end, we had to increase the current limit on the positive rail sorensen (+18 V) from ~ 7 to > 8.0 Amps to feed all the instruments. We also increased the negative (-18 V) current limit proportionally.

We think we are ready for all the new SOS on this side electronics-wise.


Photos: https://photos.app.goo.gl/GviuqLQviSPo1M3G6

ELOG V3.1.3-